Fonts from the Xy-Pic package and Layout of 10 point fonts

upper arrow tips for XY-pic  xyatip10
lower squiggles/quarter circles for XY-pic xybsql10
lower arrow tips for XY-pic  xybtip10
1/8 circles with varying radii for XY-pic xycirc10
dashes for XY-pic  xydash10
line segments for XY mode  xyline10
miscellaneous characters for XY mode  xymisc10

Layout of xyatip10 (upper arrow tips for XY-pic at 10 point)

’0  ’1  ’2  ’3  ’4  ’5  ’6  ’7
[figure: ’0x]

˝ 0x

[figure: ’1x]
[figure: ’2x]

˝ 1x

[figure: ’3x]
[figure: ’4x]

˝ 2x

[figure: ’5x]
[figure: ’6x]

˝ 3x

[figure: ’7x]
[figure: ’8x]

˝ 4x

[figure: ’9x]
[figure: ’10x]

˝ 5x

[figure: ’11x]
[figure: ’12x]

˝ 6x

[figure: ’13x]
[figure: ’14x]

˝ 7x

[figure: ’15x]
˝ 8  ˝ 9  ˝ A  ˝ B  ˝ C  ˝ D  ˝ E  ˝ F

Layout of xybsql10 (lower squiggles/quarter circles for XY-pic at 10 point)

’0  ’1  ’2  ’3  ’4  ’5  ’6  ’7
[figure: ’0x]

˝ 0x

[figure: ’1x]
[figure: ’2x]

˝ 1x

[figure: ’3x]
[figure: ’4x]

˝ 2x

[figure: ’5x]
[figure: ’6x]

˝ 3x

[figure: ’7x]
[figure: ’8x]

˝ 4x

[figure: ’9x]
[figure: ’10x]

˝ 5x

[figure: ’11x]
[figure: ’12x]

˝ 6x

[figure: ’13x]
[figure: ’14x]

˝ 7x

[figure: ’15x]
˝ 8  ˝ 9  ˝ A  ˝ B  ˝ C  ˝ D  ˝ E  ˝ F

Layout of xybtip10 (lower arrow tips for XY-pic at 10 point)

’0  ’1  ’2  ’3  ’4  ’5  ’6  ’7
[figure: ’0x]

˝ 0x

[figure: ’1x]
[figure: ’2x]

˝ 1x

[figure: ’3x]
[figure: ’4x]

˝ 2x

[figure: ’5x]
[figure: ’6x]

˝ 3x

[figure: ’7x]
[figure: ’8x]

˝ 4x

[figure: ’9x]
[figure: ’10x]

˝ 5x

[figure: ’11x]
[figure: ’12x]

˝ 6x

[figure: ’13x]
[figure: ’14x]

˝ 7x

[figure: ’15x]
˝ 8  ˝ 9  ˝ A  ˝ B  ˝ C  ˝ D  ˝ E  ˝ F

Layout of xycirc10 (1/8 circles with varying radii for XY-pic at 10 point)

’0  ’1  ’2  ’3  ’4  ’5  ’6  ’7
[figure: ’0x]

˝ 0x

[figure: ’1x]
[figure: ’2x]

˝ 1x

[figure: ’3x]
[figure: ’4x]

˝ 2x

[figure: ’5x]
[figure: ’6x]

˝ 3x

[figure: ’7x]
[figure: ’8x]

˝ 4x

[figure: ’9x]

[figure: ’10x]

˝ 5x

[figure: ’11x]

[figure: ’12x]

˝ 6x

[figure: ’13x]

[figure: ’14x]

˝ 7x

[figure: ’15x]

˝ 8  ˝ 9  ˝ A  ˝ B  ˝ C  ˝ D  ˝ E  ˝ F

Layout of xydash10 (dashes for XY-pic at 10 point)

’0  ’1  ’2  ’3  ’4  ’5  ’6  ’7
[figure: ’0x]

˝ 0x

[figure: ’1x]
[figure: ’2x]

˝ 1x

[figure: ’3x]
[figure: ’4x]

˝ 2x

[figure: ’5x]
[figure: ’6x]

˝ 3x

[figure: ’7x]
[figure: ’8x]

˝ 4x

[figure: ’9x]
[figure: ’10x]

˝ 5x

[figure: ’11x]
[figure: ’12x]

˝ 6x

[figure: ’13x]
[figure: ’14x]

˝ 7x

[figure: ’15x]
˝ 8  ˝ 9  ˝ A  ˝ B  ˝ C  ˝ D  ˝ E  ˝ F

Layout of xyline10 (line segments for XY mode at 10 point)

’0  ’1  ’2  ’3  ’4  ’5  ’6  ’7
[figure: ’0x]

˝ 0x

[figure: ’1x]
[figure: ’2x]

˝ 1x

[figure: ’3x]
[figure: ’4x]

˝ 2x

[figure: ’5x]
[figure: ’6x]

˝ 3x

[figure: ’7x]
[figure: ’8x]

˝ 4x

[figure: ’9x]
[figure: ’10x]

˝ 5x

[figure: ’11x]
[figure: ’12x]

˝ 6x

[figure: ’13x]
[figure: ’14x]

˝ 7x

[figure: ’15x]
˝ 8  ˝ 9  ˝ A  ˝ B  ˝ C  ˝ D  ˝ E  ˝ F

Layout of xymisc10 (miscellaneous characters for XY mode at 10 point)

’0  ’1  ’2  ’3
[figure: ’0x]

˝ 0x

[figure: ’1x]
[figure: ’2x]

˝ 1x

[figure: ’3x]
[figure: ’4x]

˝ 2x

[figure: ’5x]

[figure: ’6x]

˝ 3x

[figure: ’7x]

[figure: ’8x]

˝ 4x

[figure: ’9x]

[figure: ’10x]

˝ 5x

[figure: ’11x]

[figure: ’12x]

˝ 6x

’13x

’14x

˝ 7x

’15x

˝ 8  ˝ 9  ˝ A  ˝ B

Layout of xymisc10 (miscellaneous characters for XY mode at 10 point)

’4  ’5  ’6  ’7
[figure: ’0x]

˝ 0x

[figure: ’1x]
[figure: ’2x]

˝ 1x

[figure: ’3x]

[figure: ’4x]

˝ 2x

[figure: ’5x]

[figure: ’6x]

˝ 3x

[figure: ’7x]

[figure: ’8x]

˝ 4x

[figure: ’9x]

[figure: ’10x]

˝ 5x

[figure: ’11x]

[figure: ’12x]

˝ 6x

’13x

’14x

˝ 7x

’15x

˝ C  ˝ D  ˝ E  ˝ F